OpenStudy (anonymous):

Graph the system of linear inequalities. y ≤ -2x + 2 y ≥ x – 6  how do you solve a question like this?

OpenStudy (anonymous):

well first graph it as if it were normal y=mx+b

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it is this i shaded it in :)

OpenStudy (anonymous):

well once you graph a normal y=mx+b form just look at ^ this if it is < or <= shade below if it is => or > shade above. ooh if it is > < then use a doted line not a solid line.

OpenStudy (anonymous):

what are the steps to put it into graph form?

OpenStudy (binary3i):

first draw the lines on the graph. then yous the values of the x and get y. look for the valyes of x that fit in the constrains.

OpenStudy (anonymous):

use y=mx + b y ≤ -2x + 2 y= -2 x + 2 the line crosses the y axix at 2 from there it rises -2 and runs 1 -2/1 rise over run :)